Circular No. 8419

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ey IN UGC 11816 Further to IAUC 8418, O. Trondal and M. Schwartz report their discov-

ery of an apparent supernova (mag about 16.0) on unfiltered CCD images taken with the 0.81-m Tenagra II telescope on Oct. 14.2 and 16.1 UT. SN 2004ey is located at α = 21^{h}49^{m}07^{s}.81, δ = +0^{o}26′39′′.2 (equinox 2000.0), which is 8′′.0 east and 12′′.8 south of the nucleus of UGC 11816. SN 2004ey is not present on Digitized Sky Survey images from 1954 and 1986 (limiting mag ∼ 21).

SUPERNOVA 2004ez IN NGC 3430 S. Nakano, Sumoto, Japan, reports the discovery by K. Itagaki (Teppo-

cho, Yamagata, Japan, 0.60-m reflector) of an apparent supernova (red mag 17.3) on unfiltered CCD images taken on Oct. 15.82 and 16.74 UT, with a prediscovery frame taken on Sept. 30.83 showing the new object at mag 17.2. SN 2004ez is located at α = 10^{h}52^{m}15^{s}.11, δ = +32^{o}57′20′′.1 (equinox 2000.0), which is 46′′ east and 18′′.7 north of the center of NGC 3430. Nothing is visible at this location on Itagaki’s survey frames from 2000 Nov. 5 and 2001 Mar. 24 (limiting mag 18.5) or on Digitized Sky Survey images. K. Kadota (Ageo, Japan, 0.25-m f/5 reflector + unfiltered CCD) confirmed the presence of the new object at mag 17.2 on Oct. 16.79.

COMETS C/2004 P6, P7, Q3–Q6, R4–R7, AND T2 (SOHO) All of the following SOHO comets are Kreutz sungrazers except C/2004

T2 (Meyer group); continuation to IAUC 8398 (KB = K. Battams):

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 P6 Aug. 13.113 9^{h}10.^{m}2 +13^{o}24′ C3  JD 2004-T66
C/2004 P7  15.221 9 13.0 +12 19 C3  TH 2004-T66
C/2004 Q3  19.888 9 33.0 +10 50 C3  TH 2004-T66
C/2004 Q4  20.929 9 37.5 +10 35 C3  JS 2004-T66
C/2004 Q5  25.488 9 59.6 + 9 37 C3  XL 2004-T66
C/2004 Q6  26.013 9 57.9 + 9 15 C3  TH 2004-T66
C/2004 R4 Sept. 4.596 10 24.5 + 5 31 C3/2 HS 2004-T67
C/2004 R5  4.971 10 38.9 + 6 04 C3  JS 2004-T67
C/2004 R6  5.679 10 41.1 + 5 47 C3  HS 2004-T67
C/2004 R7  5.971 10 36.5 + 5 21 C3/2 KB 2004-T67
C/2004 T2 Oct. 8.998 12 55.5 − 4 31 C2  RK 2004-T76
2004 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green